Susie’s Radio Store has expected sales of $2,000,000 and plans to produce $2,500,000 worth of goods. To meet production, the business purchases $100,000 of new equipment. Actual sales for the year wind up being $1,800,000. For Susie’s Radio Store actual investment equals ______ and planned investment equals ______.

A. $600,000; $200,000

B. $800,000; $600,000

C. $500,000; $800,000

D. $650,000; $500,000
Answer #1

Option (A).

Actual investment = 2,500,000 - 2,000,000 + 100,000 = 600,000

Unplanned inventory accumulation = 2,000,000 - 1,800,000 = 200,000